OpenStudy (anonymous):

−3(x + 3) = −3(x + 1) − 5

OpenStudy (anonymous):

solve for x?

OpenStudy (phi):

-3(x+3) means -3 times each term inside the parens in other words -3(x+3) can be re-written as -3*x + -3*3 so your problem can be written as -3*x + -3*3 = -3(x+1) - 5 can you do the same thing for the right side ?

OpenStudy (anonymous):

it would end up being -3x -9 = -3x - 8 which doesn't make sense unless x is infinity

OpenStudy (phi):

yes, it looks like there is no solution for this "equation"